OpenStudy (anonymous):

what do you do when your given a point and a given line? point is (-8,3) line is 2x+3y=5 im trying to find a parallel line and a perpendicular line

OpenStudy (anonymous):

using that same point?

OpenStudy (lanhikari22):

trying to find a parallel line and a perpendicular line?, ok first solve the function to get any other point, then calculate delta x-component, delta y-component of both points, creating a vector. <x2-x1,y2-y1> so let's say you end up with vector A <x,y> find its magnitude |A| by using the phythagorean theory |A| = squareroot(x^2 + y^2), now that you've got |A| just find the angle that created our nice vector. so tan phi = y/x, meaning phi = tan^-1 y/x. now that you've got both the magnitude and the angle, just create the x- y- components. of your very very nice vector. x = |A| cos phi y = |A| sin phi x- and y- are both perpendicular to each other. well that's just what i figured out, maybe that's not what you need. maybe...try to specify a bit more? sorry couldn't help.

OpenStudy (aum):

The given line is 2x+3y=5 Put this equation in slope-intercept form and find the slope of the given line. Any line PARALLEL to the above line will have the SAME slope. We want the parallel line to pass through (-8,3). Use the point-slope formula to find the equation of the parallel line. Put it in the standard form. If a given line has a slope 'm', then any line perpendicular to the given line will have the slope -1/m. We want the perpendicular line to pass through (-8,3). Use the point-slope formula to find the equation of the perpendicular line. Put it in the standard form.